there are 216 people in a theatre. 1/3 of the male audience and 3/4 of the female audience are children. There is an equal number of male and female audience. How many children are there in the theatre?
Answer by mananth(16946)

there are 216 people in a theatre. 1/3 of the male audience and 3/4 of the female audience are children. There is an equal number of male and female audience. How many children are there in the theatre?
216 people are there in the auditorium.
let the male and female audience be x
x/3 +3x/4 are children
x+x + x/3 + 3x/4 =216
2x +x/3 + 3x/4 = 216
LCM =12
24x +4x+9x = 12*216
37x = 2592
x= 2592/37
x=70
Female , Men = 70
children = 76